OpenStudy (anonymous):

What is the solution set for |–x + 2| = 5?

OpenStudy (zarkon):

solve (-x+2)=5 and -(-x+2)=5 for x

OpenStudy (anonymous):

how @Zarkon

OpenStudy (zarkon):

with basic algebra

OpenStudy (anonymous):

okay i got -3 how do i get the other set ? @zarkon

OpenStudy (anonymous):

okay i got -3 for one is that correct ?

OpenStudy (anonymous):

a positive 7

OpenStudy (anonymous):

can i ask for help in another question ? @hero

OpenStudy (anonymous):

y varies inversely as twice x. When x = 4, y = 10. Find y when x = 8

OpenStudy (anonymous):

y = 5

OpenStudy (anonymous):

what about this question What is the equation of a vertical line passing through the point (–4, 7)?

OpenStudy (anonymous):

god @hero i love you so much for helping me

OpenStudy (anonymous):

-4 ?

OpenStudy (anonymous):

To which graph does the point (–2, –1) belong? thank you just a couple more its juts cus i only understand you

OpenStudy (anonymous):

okay thank you and this one lol im so sorry love i just need the help What is the equation in point-slope form of the line passing through (–1, 3) and (1, 7)?

OpenStudy (anonymous):

is it y – 3 = 4(x + 1)

hero (hero):

Why would that be it? If you simplify that you get y = 4x + 7, which is not the proper form. I guess you didn't understand the steps I provided, which is sad because I explained every step.

OpenStudy (anonymous):

so its y – 7 = 4(x – 1)

OpenStudy (anonymous):

i did understand i must have missed something

OpenStudy (anonymous):

i already got the answer i need help on a new question

OpenStudy (anonymous):

y varies jointly with p and q. When p = 3 and q = 5, y = 60. Find y when p = 2 and q = 6.

OpenStudy (anonymous):

@Hero you their

OpenStudy (anonymous):

okay i 3=2(-1) = 5 which is 3=-2=5 which is 3